I also need help in this one Please, It's very urgent.
Suppose a life insurance policy costs $16 for the first unit of coverage and then $4 for each additional unit of coverage. Let C(x) be the cost for insurance of x units of coverage. What will 10 units of coverage cost?
Thanks

Answer by Earlsdon(6294)

C(x) = 16 + 4(x-1) "$16 for the first unit plus $4 for each additional unit after the first" This function can be simplified.
C(x) = 16+4x-4
C(x) = 12+4x To find the cost of 10 units. substitute x = 10.
C(10) = 12+4(10)
C(10) = 12+40
C(10) = 52
10 units of coverage will cost $52